# Wave-U-Mamba: An End-To-End Framework For High-Quality And Efficient Speech Super Resolution

### v2

@@ Abstract
Speech Super-Resolution (SSR) is a task of enhancing low-resolution speech signals by restoring missing high-frequency components. Conventional approaches typically reconstruct log-mel features, followed by a vocoder that generates high-resolution speech in the waveform domain. However, as log-mel features lack phase information, this can result in performance degradation during the reconstruction phase. Motivated by recent advances with Selective State Spaces Models (SSMs), we propose a method, referred to as Wave-U-Mamba that directly performs SSR in time domain. In our comparative study, including models such as WSRGlow, NU-Wave 2, and AudioSR, Wave-U-Mamba demonstrates superior performance, achieving the lowest Log-Spectral Distance (LSD) across various low-resolution sampling rates, ranging from 8 kHz to 24 kHz. Additionally, subjective human evaluations, scored using Mean Opinion Score (MOS) reveal that our method produces SSR with natural and human-like quality. Furthermore, Wave-U-Mamba achieves these results while generating high-resolution speech over nine times faster than baseline models on a single A100 GPU, with parameter sizes less than 2% of those in the baseline models.
Index Terms—  Speech Super-Resolution, U-Net, Mamba, Generative Adversarial Networks
@@ I Introduction
SSR is a generative task that leverages the power of modern generation backbones such as diffusion probabilistic models [1, 2, 3], Generative Adversarial Networks (GAN) [4, 5], Generative Flow [6]. One way to approach SSR involves generating HR waveform directly from LR waveform [7, 8]. However, given that speech waveforms are characterized by significantly longer sequence lengths compared to mel spectrogram representations, conventional works first map the LR mel spectrogram to HR mel spectrogram and then project the recovered mel spectrogram onto a waveform through a well-defined vocoder [3, 4]. This two-step approach offers the advantage of decomposing the SSR task into simpler modules, akin to the traditional two-stage modeling employed in tasks such as Text-to-Speech (TTS) [9, 10, 11].
Despite the promising results achieved by previous works, certain limitations remain. The mel spectrogram, being a feature-reduced representation, inherently discards phase information during the conversion process. This introduces an additional challenge of implicit phase reconstruction during the SSR process. Previous works incorporate a pretrained vocoder to solve the problem so that the SSR frameworks can only aim at recovering the high-frequency components. However, this approach prevents end-to-end training of the model and is inefficient, as the phase information could be preserved if the waveform were directly utilized.
Considering these factors, it is hypothesized that a model capable of directly generating HR waveforms from LR waveforms in an efficient and effective manner would offer a more optimal solution. In this paper, motivated by the success of selective SSMs [12] to extract long-term dependent information both effectively and efficiently [13, 14], we propose Wave-U-Mamba to solve SSR in waveform domain. Wave-U-Mamba is a GAN-based framework that takes LR speech as input, with sampling rates ranging from 4 kHz to 24 kHz, and up-samples it to HR speech with a target sampling rate of 48 kHz. Owing to the inherent capabilities of the Mamba architecture to efficiently process long sequences, coupled with its U-shaped model architecture that allows direct projection from LR to HR waveforms, the proposed model can generate HR waveforms much more rapidly than existing baseline models.
@@ II Related Works
A Selective State Spaces Models
@@
B Time Domain Speech Processing
Although time-frequency representations such as mel spectrogram, Mel-Frequency Cepstral Coefficients (MFCC) [19], and Power-Normalized Cepstral Coefficients (PNCC) [20] are traditionally employed to address speech-related tasks, time domain speech processing has been studied in various tasks including source separation [21, 22], representation learning [23], multimodal processing [24], speech generation [25, 26], and speech super-resolution [7].
Fig. 1: The architecture of DownsampleBlock (Left), Wave-U-Mamba Generator (Middle), and UpsampleBlock (Right).
@@ III Wave-U-Mamba
Wave-U-Mamba is a GAN-based generative model that comprises a single generator and two discriminators. Our generator is a U-Net based architecture that takes raw waveform of LR speech as input and outputs HR speech waveform. A residual connection is employed every block to enhance stability. Especially the residual block at the final upsampling layer enables the model to focus on estimating the upper-frequency components, given that the lower-frequency features are already captured in the LR waveform. Fig. 1 illustrates the overall architecture of the Wave-U-Mamba generator. Building on the previous work [27], we utilize Multi-Period Discriminator (MPD) and Multi-Scale Discriminator (MSD), each comprising groups of sub-discriminators to perform adversarial training.
A Task Definition
@@
A LR speech signal has shape of $X_{L}=[x_{1,...,s\times L}]$. $s$ represents signal length in seconds, and $L$ represents the sampling rate of the signal, where the upper-frequency limit is set to $L/2$ by Nyquist Theorem. A HR signal has shape of $Y_{H}=[y_{1,...,s\times H}]$. $H$ represents the sampling rate of the signal, where the upper-frequency limit is set to $H/2$. We first upsample the LR signal to match the sampling rate of the HR signal (i.e. $X_{H}=[x_{1,...,s\times H}]$) using sinc interpolation. Then we parameterize and train the function $F$ so that $Y_{H}=F(X_{H})$.
B Generator
Fig. 2: The architecture of MambaBlock (Left) and ResidualBlock (Right).
@@ B Generator
Downsampling Enhancing long-range dependency is a critical challenge in feature extraction within the waveform domain since approaching distant features is more difficult compared to mel spectrogram. Previous studies have primarily focused on modifying the hyperparameters of convolutional neural networks (CNN) to extract long-term dependencies in waveform domain [25, 23]. However, these approaches often involve deepening network layers or expanding kernel sizes, leading to the loss of valuable information or increased computational complexity. To address this issue, we propose the incorporation of Mamba as a global feature extractor, due to its capability to efficiently and effectively learn long-range patterns. Specifically, we introduce the MambaBlock that integrates Mamba with Layer Normalization (LN) [28] and residual connection. We set $M$, the number of MambaBlocks to 2.
@@
Mamba-based models are recognized for their difficulty in training, largely due to the sharp or occasionally non-convex nature of their loss landscapes [29]. To mitigate this challenge, and motivated by [30], we design a 1D version of stem embedding as an operator to map input representations into a high-dimensional latent space, facilitating more stable training of the Mamba model. A StemEmbedBlock is made of 1D convolution with a kernel size of 4 and stride of 1, followed by LN, LeakyReLU, and residual connections. In practice, we use $N=2$ blocks as a function to double the feature dimension, which further stabilizes the training process. The hidden dimension at the bottleneck layer is set to 256, which is comparatively lower than that of other U-Net based SSR architectures [4, 7]. By constraining the hidden dimension size, the model effectively limits the propagation of extraneous patterns from the interpolated LR signal $X_{H}$, such as those generated by sinc interpolation, ensuring that only relevant and meaningful features are captured.
@@
Traditional approaches such as strided CNNs with padding for sequence length reduction, can introduce boundary artifacts and disrupt the continuous nature of the signal [22]. To address the problem, we use average pooling to halve the sequence length. In this way, we compress features that add to the naturalness of the generated HR signal. The bottleneck block follows the same structure as the DownsampleBlock but without pooling operation.
@@
Upsampling We employ transposed convolution as an up-sampling operator, in which kernel sizes are set as a multiple of strides to remove checkerboard artifacts common in transposed convolution [31]. MambaBlocks with the same architecture in down-sampling are utilized to capture global features. ResidualBlocks are put after Mambablocks to further improve the connectivity of representations. Specifically, every ResidualBlock has kernel sizes of 3 with dilation rates set to $d=[1,3]$. The architecture of ResidualBlock can be found in Fig. 2. The final UpsampleBlock consists of a single convolutional layer, followed by a tanh activation to predict the HR signal. All convolutional layers are weight-normalized [32].
C Discriminator
@@
Following early works on incorporating GAN for vocoding [27, 33], we use Multi-Period Discriminator (MPD) and Multi-Scale Discriminator (MSD) as two discriminators to conduct adversarial training of Wave-U-Mamba. MPD first reshapes speech signal into 2D data and then applies 2D convolution to capture periodic patterns in the signal [27]. MPD consists of 5 sub-discriminators with different periods to reshape speech signals. MSD is a 1D convolution-based discriminator that operates on smoothed waveform [33]. MSD consists of 3 sub-discriminators that operate on raw waveform and two down-sampled waveforms. By combining two modules, our generator is expected to capture both periodic patterns and continuous nature of speech signals, rendering natural and high-quality generation of HR speech. Given that generators used in SSR tasks are more complex than generators used in vocoding, we deepen the structure of both MPD and MSD by adding more building blocks to balance training between generator and discriminator.
D Training Objectives
@@
Mel Spectrogram Loss For the natural and realistic SSR, it is crucial that perceptual quality be measured thoroughly. We therefore use L1-loss of the mel spectrogram of the predicted signal and the ground truth signal, where $\psi$ refers to the mel transformation. In equation 1, $Y$ and $\hat{Y}$ refer to the ground truth HR speech waveform and the predicted waveform, respectively.
$L_{\text{mel}}=\mathbb{E}[||\psi(Y)-\psi(\hat{Y})||_{1}]$ (1)
@@
Multi-Resolution Short-Time Fourier Transform (STFT) Loss Since SSR is a task that aims to reconstruct high-frequency components, we incorporate Multi-Resolution Short-Time Fourier Transform (STFT) Loss [34] as another loss term for the model to learn detailed characteristics of high-frequency bands in the spectrogram. Multi-Resolution STFT Loss is the sum of the STFT losses with different parameters (i.e.window size, hop length). Each STFT loss $L_{\mathrm{s}}$ is the sum of spectral convergence loss and log STFT magnitude loss. It also helps prevent learning the patterns of specific hyper-parameters used in STFT.
$L_{\text{STFT}}(Y,\hat{Y})=\mathbb{E}[L_{\text{s}}(Y,\hat{Y})]$ (2)
@@
$L_{\text{GAN}}(D;G)$ $=\mathbb{E}\left[(D(x)-1)^{2}+\left(D(G(s))\right)^{2}\right]$ (3)
$L_{\text{GAN}}(G;D)$ $=\mathbb{E}\left[(D(G(s))-1)^{2}\right]$
@@
We set $\lambda_{\text{mel}}=45$ and $\lambda_{\text{STFT}}=10$. The coefficients were set to balance the values among loss terms. We did not use L1 Loss of the two signals because minimizing the distance between two waveforms does not direct to better perceptual quality nor the recovery of the high-frequency components, both intuitively and empirically.
@@ IV Experimental Results
In this section, we describe the training settings and evaluation criteria used in our experiments. Then we visualize some sample mel spectrograms to help understand our results, which is followed by both objective and subjective score comparison of our models with other baselines, including WSRGlow [6], Nu-Wave 2 [1] and AudioSR [3]. Furthermore, we show the results of ablation studies to demonstrate characteristics of the design choices. Lastly, we show the efficiency measure of our model compared with other models.
A Training
@@
VCTK [35] is used as a dataset both for training and validation. The data is first cut into VCTK-train and VCTK-test, following the same data-splitting scheme from the previous work [7] then sampled at 48 kHz. We chunk about 0.7 seconds of data to use for training. We randomly set the cutoff frequency between 2 kHz and 12 kHz and then pass the data through an order-8 Chebyshev filter. The signal is down-sampled to a target sampling rate and then up-sampled back to 48 kHz to remove the high-frequency components completely. We normalize the signal onto $[-1,1]$ in training because our model has tanh activation on the last layer to predict the HR signal.
Every model is trained for 25 epochs with early stopping to prevent overfitting. AdamW is used as an optimizer with $\beta_{1}=0.6$ and $\beta_{2}=0.99$. The learning rate is initially set to $42e-$ 4 with warm-up and weight decay. Experiments were conducted under 2 RTX A6000 GPUs with batch size set to 64, which takes roughly 8 hours to train our model.
B Evaluation
@@
The primary evaluation metric for objective evaluation is Log-Spectral Distance (LSD), illustrated in Equation 5. $Y(f,t)$ and $\hat{Y}(f,t)$ refer to the magnitude spectrogram of the ground truth signal and the predicted signal. LSD can measure how well the signal has recovered the frequency components.
$\vspace{-1.0em}\mathrm{LSD}(Y,\hat{Y})=\frac{1}{T}\sum_{t=1}^{T}\sqrt{\frac{1}{F}\sum_{f=1}^{F}\log_{10}\left(\frac{Y(f,t)^{2}}{\hat{Y}(f,t)^{2}}\right)^{2}}$ (5)
@@
To evaluate the perceptual quality of the generated speech, we implement 5-scale Mean Opinion Score (MOS) tests on samples from WSRGlow [6], Wave-U-Mamba, and AudioSR [3] then provide MOS scores. 45 raters tested on 5 speech samples randomly selected from the VCTK-test. All samples were anonymized and volume-normalized to guarantee the fairness of the testing procedure.
@@
Fig. 3: Enhanced mel spectrograms using Wave-U-Mamba at different training phases. 3a and 3b represent Low-Resolution and High-Resolution mel spectrogram of the ground truth data. 3c and 3d represent recovered mel spectrogram from epoch 2 and 9, respectively.
C Results

### v3

@@ Abstract
Speech Super-Resolution (SSR) is a task of enhancing low-resolution speech signals by restoring missing high-frequency components. Conventional approaches typically reconstruct log-mel features, followed by a vocoder that generates high-resolution speech in the waveform domain. However, as mel features lack phase information, this can result in performance degradation during the reconstruction phase. Motivated by recent advances with Selective State Spaces Models (SSMs), we propose a method, referred to as Wave-U-Mamba that directly performs SSR in time domain. In our comparative study, including models such as WSRGlow, NU-Wave 2, and AudioSR, Wave-U-Mamba demonstrates superior performance, achieving the lowest Log-Spectral Distance (LSD) across various low-resolution sampling rates, ranging from 8 to 24 kHz. Additionally, subjective human evaluations, scored using Mean Opinion Score (MOS) reveal that our method produces SSR with natural and human-like quality. Furthermore, Wave-U-Mamba achieves these results while generating high-resolution speech over nine times faster than baseline models on a single A100 GPU, with parameter sizes less than 2% of those in the baseline models.
Index Terms:
Speech Super-Resolution, U-Net, Mamba, Generative Adversarial Networks
@@ I Introduction
SSR is a generative task that leverages the power of modern generation backbones such as diffusion probabilistic models [1, 2, 3], Generative Adversarial Networks (GAN) [4, 5], Generative Flow [6]. One way to approach SSR involves generating HR waveform directly from LR waveform [7, 8, 9]. However, given that speech waveforms are characterized by significantly longer sequence lengths compared to mel spectrogram representations, conventional works first map the LR mel spectrogram to HR mel spectrogram and then project the recovered mel spectrogram onto a waveform through a well-defined vocoder [3, 4]. This two-step approach offers the advantage of decomposing the SSR task into simpler modules, akin to the traditional two-stage modeling employed in tasks such as Text-to-Speech (TTS) [10, 11, 12].
Despite the promising results achieved by two-step approaches, certain limitations remain. The mel spectrogram, being a feature-reduced representation, inherently discards phase information during the conversion process. This introduces an additional challenge of implicit phase reconstruction during the SSR process. Previous works incorporate a pretrained vocoder to solve the problem so that SSR frameworks can only aim at recovering the high-frequency magnitudes. However, this approach often does not render fully end-to-end training of the model [4, 3, 11] and is inefficient, as phase information could be preserved if the waveform were directly utilized.
Considering these factors, it is hypothesized that a model capable of directly generating HR waveforms from LR waveforms in an efficient and effective manner would offer a more optimal solution. In this paper, motivated by the success of selective SSMs [13] to extract long-term dependent information both effectively and efficiently [14, 15], we propose Wave-U-Mamba to solve SSR in waveform domain. Wave-U-Mamba is a GAN-based framework that takes LR speech as input, with sampling rates ranging from 4 to 24 kHz, and upsamples it to HR speech with a target sampling rate of 48 kHz. Owing to the inherent capabilities of the Mamba architecture to efficiently process long sequences, coupled with its U-shaped model architecture that allows direct projection from LR to HR waveforms, the proposed model can generate HR waveforms much more rapidly than existing baseline models. Audio samples are made available online. [^1] Listenable demos are available at infected4098.github.io/waveumambademo.
@@ II Related Works
II-A Selective State Spaces Models
@@
II-B Time Domain Speech Processing
Although time-frequency representations such as mel spectrogram are traditionally employed to address speech-related tasks, time domain speech processing has been studied in various tasks including source separation [20, 21], representation learning [22], multimodal processing [23], speech generation [24, 25], and speech super-resolution [7, 8, 9].
Figure 1: The architecture of DownsampleBlock (Left), Wave-U-Mamba Generator (Middle), and UpsampleBlock (Right).
@@ III Wave-U-Mamba
Wave-U-Mamba is a GAN-based generative model that comprises a single generator and two discriminators. Our generator is a U-Net based architecture that takes the raw waveform of LR speech as input and outputs HR speech waveform. A residual connection is employed every block to enhance stability. Especially, the residual block at the final upsampling layer enables the model to focus on estimating the upper-frequency components, given that the lower-frequency features are already captured in the LR waveform. Fig. 1 illustrates the overall architecture of the Wave-U-Mamba generator. Building on the previous work [26], we utilize Multi-Period Discriminator (MPD) and Multi-Scale Discriminator (MSD), each comprising groups of sub-discriminators to perform adversarial training.
III-A Task Definition
@@
A LR speech signal has shape of $X_{L}=[x_{1,...,s\times L}]$. $s$ represents signal length in seconds, and $L$ represents the sampling rate of the signal, where the upper-frequency limit is set to $L/2$ by Nyquist Theorem. A HR signal has shape of $Y_{H}=[y_{1,...,s\times H}]$. $H$ represents the sampling rate of the signal, where the upper-frequency limit is set to $H/2$. We first upsample the LR signal to match the sampling rate of the HR signal (i.e. $X_{H}=[x_{1,...,s\times H}]$) using Fast Fourier Transform (FFT) interpolation. Then we parameterize and train the generative function $G$ so that $Y_{H}=G(X_{H})$.
III-B Generator
Figure 2: The architecture of MambaBlock (Left) and ResidualBlock (Right).
@@ III-B Generator
Downsampling Enhancing long-range dependency is a critical challenge in feature extraction within the waveform domain. Approaching distant features in waveform domain is harder than time-frequency domain due to the comparatively longer duration that a single frame in spectrogram spans. Previous studies have primarily focused on modifying the hyperparameters of convolutional neural networks (CNN) to extract long-term dependencies in waveform domain [24, 22]. However, these approaches often involve deepening network layers or expanding kernel sizes, leading to the loss of valuable information or increased computational complexity. To address this issue, we propose the incorporation of Mamba as a global feature extractor, due to its capability to efficiently and effectively learn long-range patterns. Specifically, we introduce MambaBlock that integrates Mamba with Layer Normalization (LN) [27] and residual connection. We set $M$, the number of MambaBlocks to 2.
@@
Mamba-based models are recognized for their difficulty in training, largely due to the relatively sharp or non-convex nature of their loss landscapes [28]. To mitigate this challenge, and motivated by [29], we design a 1D version of stem embedding as an operator to map input representations into a high-dimensional latent space, facilitating more stable training of the Mamba model. A StemEmbedBlock is made of 1D convolution with a kernel size of 4 and a stride of 1, followed by LN, LeakyReLU, and residual connection. In practice, we set $N$, the number of StemEmbedBlocks to 2 to double the feature dimension, which further stabilizes the training process. The hidden dimension at the bottleneck layer is set to 256, which is comparatively lower than that of other U-Net based SSR architectures [4, 7]. By constraining the hidden dimension size, the model effectively limits the propagation of extraneous patterns from the interpolated LR signal $X_{H}$, such as those generated by FFT interpolation, ensuring that only relevant and meaningful features are captured.
@@
As shown in previous work [21], approaches such as strided CNNs with padding for sequence length reduction can often introduce borderline artifacts and disrupt the continuous nature of the signal due to the padding operation. To address the problem, we use average pooling to halve the sequence length. In this way, we compress features that add to the continuity of the generated HR signal. The bottleneck block follows the same structure as the DownsampleBlock but without pooling operation.
@@
Upsampling We employ transposed convolution as an upsampling operator in which kernel sizes are set as a multiple of strides to reduce checkerboard artifacts common in transposed convolution [30]. MambaBlocks with the same architecture in downsampling are utilized to capture global features. ResidualBlocks are put after Mambablocks to further improve the connectivity of representations and enhance the training stability. Specifically, every ResidualBlock has kernel sizes of 3 with dilation rates set to $d=[1,3]$. The architecture of ResidualBlock can be found in Fig. 2. Unlike other UpsampleBlocks, the final UpsampleBlock consists of a single convolutional layer, followed by a tanh activation to predict the residual HR signal. All convolutional layers are weight-normalized [31].
III-C Discriminator
@@
Following early works on incorporating GAN for vocoding [26, 32], we use Multi-Period Discriminator (MPD) and Multi-Scale Discriminator (MSD) as two discriminators to conduct adversarial training of Wave-U-Mamba. MPD first reshapes speech signal into 2D data and then applies 2D convolution to capture periodic patterns in the signal [26]. MPD consists of 5 sub-discriminators with different reshaping periods to reshape speech signals. MSD is a 1D convolution-based discriminator that operates on smoothed waveform [32]. MSD consists of 3 sub-discriminators that operate on raw waveform and two downsampled waveforms. By combining two modules, our generator is expected to capture both periodic patterns and continuous nature of speech signals, rendering natural and high-quality generation of HR speech. Given that the discriminators in SSR tasks deal with comparatively fine-grained HR speech signal than those used in common vocoding tasks, we deepen the structure of both MPD and MSD by adding more building blocks.
III-D Training Objectives
@@
Mel Spectrogram Loss For the natural and realistic SSR, it is crucial that perceptual quality be measured thoroughly. We therefore use L1-loss of the mel spectrogram of the predicted signal and the ground truth signal, where $\psi$ refers to the mel transformation. In equation 1, $y$ and $x$ refer to the ground truth HR and the LR waveform, respectively. A function $G$ is our generator Wave-U-Mamba that takes as input $x$ to predict the HR waveform $y$.
$L_{\text{mel}}=\mathbb{E}_{(y,x)}[||\psi(y)-\psi(G(x))||_{1}]$ (1)
@@
Multi-Resolution Short-Time Fourier Transform (STFT) Loss Since SSR is a task that aims to reconstruct high-frequency components, we incorporate Multi-Resolution Short-Time Fourier Transform (STFT) Loss [33] as another loss term for the model to learn detailed characteristics of high-frequency bands in the spectrogram. Multi-Resolution STFT Loss is the sum of the STFT losses with different parameters such as window size or hop length, where $M$ represents the number of STFT losses. Each STFT loss $L_{\mathrm{s}}$ is the sum of spectral convergence loss and log STFT magnitude loss.
$L_{\text{STFT}}(y,G(x))$ $=\frac{1}{M}\sum^{M}_{m=1}L_{\text{s}}^{(m)}(y,G(x))$ (2)
$L_{s}(y,G(x))$ $=\mathbb{E}_{(y,x)}[L_{sc}(y,G(x))+L_{mag}(y,G(x))]$
@@
$L_{\text{GAN}}(D;G)$ $=\mathbb{E}_{(y,x)}\left[(D(y)-1)^{2}+\left(D(G(x))\right)^{2}\right]$ (3)
$L_{\text{GAN}}(G;D)$ $=\mathbb{E}_{x}\left[(D(G(x))-1)^{2}\right]$
@@
We set $\lambda_{\text{mel}}=45$ and $\lambda_{\text{STFT}}=10$. The coefficients were set to balance the overall scale of loss terms. We did not use L1 Loss of the two waveforms because minimizing the distance between two waveforms does not direct to better perceptual quality nor the recovery of the high-frequency components, as pointed out in previous literature [34, 35].
@@ IV Experimental Results
In this section, we describe the training settings and evaluation criteria used in our experiments. Then we visualize some sample mel spectrograms to help understand our results, which is followed by both objective and subjective score comparison of our models with other baselines, including WSRGlow [6], Nu-Wave 2 [1] and AudioSR [3]. Furthermore, we show the results of ablation studies to demonstrate the characteristics of the design choices. Lastly, we show the efficiency measure of our model compared with other models.
IV-A Training
@@
VCTK [36] is used as a dataset both for training and validation. The data were first cut into VCTK-train and VCTK-test, following the same data splitting scheme from the previous work [7] then were sampled at 48 kHz. We segmented about 0.7 seconds of data to use for training. We randomly set the cutoff frequency between 2 kHz and 12 kHz and then passed the data through an order-8 Chebyshev type 1 filter. The signal is downsampled to a target sampling rate and then upsampled back to 48 kHz to remove the high-frequency components completely. We normalized the signal onto $[-1,1]$ in training because our model has tanh activation on the last layer to generate the HR signal. For the calculation of the mel spectrogram, we use 80 mel filterbanks, the Hanning window with a window length of 2048 and a hop length of 512 for both the training and the evaluation. For the computation of Multi-Resolution STFT loss, we followed the default configuration of the open-source tool. [^2] https://github.com/csteinmetz1/auraloss
The model was trained for 25 epochs, during which we stopped the training when the validation loss does not decrease for three consecutive epochs. AdamW[37] is used as an optimizer with $\beta_{1}=0.6$ and $\beta_{2}=0.99$. We set the learning rate to $54e-$ 5 for both the generator and the discriminator, which linearly increases to $42e-$ 4 in epoch 5 and then decreases with exponential weight decay with $\gamma$ set to 0.999. We apply gradient clipping to both the generator and the discriminator with a maximum norm of 2 to stabilize the training process. Experiments were conducted under 2 RTX A6000 GPUs with batch size set to 64, which took approximately 8 hours to train our model.
IV-B Evaluation
@@
The primary evaluation metric for the objective evaluation is Log-Spectral Distance (LSD), illustrated in Equation 5. $Y(f,t)$ and $\hat{Y}(f,t)$ refer to the magnitude spectrogram of the ground truth signal and the predicted signal. LSD can measure how well the signal has recovered the frequency components.
$\mathrm{LSD}(Y,\hat{Y})=\frac{1}{T}\sum_{t=1}^{T}\sqrt{\frac{1}{F}\sum_{f=1}^{F}\log_{10}\left(\frac{Y(f,t)^{2}}{\hat{Y}(f,t)^{2}}\right)^{2}}$ (5)
@@
To evaluate the perceptual quality of the generated speech, we implemented 5-scale Mean Opinion Score (MOS) tests on samples from WSRGlow [6], Wave-U-Mamba, and AudioSR [3] then provided MOS scores. 45 raters were tested on 5 speech samples randomly selected from the VCTK-test to measure the overall quality of the speech. The sampling rate of 8 kHz and 12 kHz were used for simulating the LR speech. All samples were anonymized and volume-normalized to guarantee the fairness of the testing procedure.
@@
Figure 3: Enhanced mel spectrograms using Wave-U-Mamba at different training phases. 3a and 3b represent Low-Resolution and High-Resolution mel spectrogram of the ground truth speech signal. 3c and 3d represent recovered mel spectrogram from epoch 2 and 9, respectively.
IV-C Results